40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1968     Entry time: Fri Jul 10 00:59:49 2015
Author: Arjun 
Type: Misc 
Category: PD noise 
Subject: PD-noise update 

Today we were able to get the servo running which we described in the previous eLog. The servo locked beautifully and the instabilities other wise observed without the zero we specifically added was gone. There was small change that had to be made, the servo had the wrong sign which we very took care of. So, just for the sake of completeness, I am attaching the TF of the inverted and the non- inverted servo. Also, I am attaching the corrected schematic. I coudn't take a snap of how the lock was before and after we introduced the our servo( I will make sure its in the next log).

A image of the servo we built






Addititonally, we assembled koji style beam dumps onto our setup, but the alignment was such that we couldn't place the third beam dump for the reflection coming from the bea splitter. So we may have to realign the beam splitter and then re-align the PDs as well. Its not as tedious as it sounds and we should be done with it tomorrow.


The only thing that would now remain would be to design and implement the servo for intensity suppression, what makes it ore difficult is that we need to include a very quiet reference in the circuit and we are brainstroming over it. Also, find attached a image of the actual servo we built. Also, I have attached the LISO simulation code.

Attachment 1: cav_inv.fig  21 kB
Attachment 2: cavity_servo.fil  589 Bytes  Uploaded Fri Jul 10 02:01:33 2015  | Hide | Hide all
#---------------Frequency tracking control servo---------------

#stage-1- non-inverting gain stage,G=30
r r1 1k n1 gnd
r r2 32.8k n1 n2

op u1 lt1128 nin n1 n2

#stage-2, Low pass filter pole=10Hz, zero=500Hz
r r4 2k n2 n3
... 29 more lines ...
Attachment 3: Screenshot_2015-07-10-00-41-15.png  2.367 MB  Uploaded Fri Jul 10 02:02:31 2015  | Hide | Hide all
Attachment 4: IMG_20150708_180717.jpg  835 kB  Uploaded Fri Jul 10 02:02:31 2015  | Hide | Hide all
Attachment 5: Screenshot_2015-07-10-00-51-01.png  3.186 MB  Uploaded Fri Jul 10 02:03:23 2015  | Hide | Hide all
Attachment 6: cavity-servo_(2).pdf  32 kB  Uploaded Fri Jul 10 12:10:45 2015  | Hide | Hide all
ELOG V3.1.3-